Tohban Report 2018-06-20
Tohban Reports | |
---|---|
Start Date: | 13 June 2018 |
End Date: | 20 June 2018 |
Tohban: | Mitsuo Oka |
Tohban email: | moka@berkeley.edu |
Next Tohban: | Pascal |
List all reports |
Solar Activity
The solar activity increased gradually from A-level to B-level, and there are now three active regions on disk. However, we still expect the sun to remain quiet for another week.
No RHESSI flares to count.
Memory Management
Detectors are off, SSR fill around at 60%
Spacecraft Status
Detectors are off, but system is not cooling due to the receiver lock and cryocooler problems. See
http://sprg.ssl.berkeley.edu/~ayshih/soh/latest_temperatures.pdf http://sprg.ssl.berkeley.edu/~ayshih/soh/latest_temperatures_3days.pdf http://sprg.ssl.berkeley.edu/~ayshih/soh/latest_temperatures_2weeks.pdf
for summary plots of recent temperatures, efficiencies and accelerometer values. Cold tips are 153 and 152 K, cold plates are at 167 and 166 K.
Data Gaps
Data gaps of more than 300s between 2018-06-12 and 2018-06-19 are as follows:
GAPS IN APP_ID = 1 (VC1-SOH) WITH PACKET RATE LT 30
GAP START TIME GAP END TIME GAP (SEC) 2018-06-12T00:00:00.000 -- 2018-06-12T17:05:00.000 61500.000 2018-06-16T12:25:00.000 -- 2018-06-16T12:40:00.000 900.00000 2018-06-17T11:35:00.000 -- 2018-06-17T12:50:00.000 4500.0000 2018-06-17T12:55:00.000 -- 2018-06-17T13:25:00.000 1800.0000 2018-06-18T11:05:00.000 -- 2018-06-18T14:15:00.000 11400.000 2018-06-19T10:40:00.000 -- 2018-06-19T14:30:00.000 13800.000 2018-06-19T21:05:00.000 -- 2018-06-20T00:00:00.000 10500.000
GAPS IN APP_ID = 154 (VC1-PMTRAS) WITH PACKET RATE LT 4
GAP START TIME GAP END TIME GAP (SEC) 2018-06-12T00:00:00.000 -- 2018-06-13T00:00:00.000 86400.000 2018-06-13T00:00:00.000 -- 2018-06-13T21:40:00.000 78000.000 2018-06-16T14:10:00.000 -- 2018-06-16T14:45:00.000 2100.0000 2018-06-16T15:00:00.000 -- 2018-06-16T15:15:00.000 900.00000 2018-06-17T11:35:00.000 -- 2018-06-17T17:15:00.000 20400.000 2018-06-17T17:50:00.000 -- 2018-06-17T18:30:00.000 2400.0000 2018-06-17T19:55:00.000 -- 2018-06-17T22:25:00.000 9000.0000 2018-06-18T11:05:00.000 -- 2018-06-18T22:25:00.000 40800.000 2018-06-19T11:55:00.000 -- 2018-06-20T00:00:00.000 43500.000
GAPS IN APP_ID = 102 (VC3-MONITOR RATES) WITH PACKET RATE LT 30
GAP START TIME GAP END TIME GAP (SEC) 2018-06-12T00:00:00.000 -- 2018-06-13T00:00:00.000 86400.000 2018-06-13T00:00:00.000 -- 2018-06-14T00:00:00.000 86400.000 2018-06-14T00:00:00.000 -- 2018-06-15T00:00:00.000 86400.000 2018-06-15T00:00:00.000 -- 2018-06-16T00:00:00.000 86400.000 2018-06-16T00:00:00.000 -- 2018-06-17T00:00:00.000 86400.000 2018-06-17T00:00:00.000 -- 2018-06-18T00:00:00.000 86400.000 2018-06-18T00:00:00.000 -- 2018-06-19T00:00:00.000 86400.000 2018-06-19T00:00:00.000 -- 2018-06-20T00:00:00.000 86400.000
(Moitor rates are off, so the full week is a gap)
Detector issues
Detectors are off.
Spacecraft Management
2018-06-13 (Wed) Communication recovered. The cryocooler had been powered off by RTS 11. Cryocooler was turned on, ramped to 33W.
2018-06-14 (Thu) The cryocooler power increased from 33 to 42 W at 2018-165-22:51:45 UTC.
2018-06-15 (Fri) Last successful command 2018-166-22:18:35 UTC.
2018-06-16 (Sat) The receiver failed to lock.
2018-06-17 (Sun) The receiver failed to lock.
2018-06-18 (Mon) The receiver failed to lock.
2018-06-19 (Tue) The cryocooler was turned off by RTS 11, but some commands reached to the spacecraft later --> Turned on the cryocooler and ramped it to 26.5W.
Decimation | Active/Vigorous |
HLAT Decimation | Rear decimation |
Night time data (fronts) | plus/minus 4 minutes from daylight |
Night time data (rears) | plus/minus 4 minutes from daylight |
Require extra passes? | No |
Requirement for moving pointer? | See above |
Attenuator operation | Nominal |
Ref. deadtime for Att.thresholds | 35.x% |
Detector problems? | See above |